OpenStudy (darkbluechocobo):

Define a second angle which will result in the same range for a projectile launched at an angle to the horizontal at 35 degrees?

OpenStudy (anonymous):

The range of a projectile is given by the equation \[R=\frac{ 2v^2\sin(\theta)\cos(\theta) }{ g }\] So you need to find another angle psi such that\[\sin(\psi)\cos(\psi)=\sin(\theta)\cos(\theta)\] I think psi and theta are called complementary angles, but i could be wrong

OpenStudy (darkbluechocobo):

*sigh* i dont know what im doing s: this is my first question like this :/

OpenStudy (anonymous):

well you've seen sines and cosines before, right ?

OpenStudy (darkbluechocobo):

Yes

OpenStudy (anonymous):

okay, your condition would be satisfied if sin(psi) = cos(theta) and vice versa

OpenStudy (anonymous):

so, what is the sine of zero degrees ?

OpenStudy (darkbluechocobo):

0?

OpenStudy (anonymous):

yes, and what angle has a cosine equal to zero ?

OpenStudy (darkbluechocobo):

45?

OpenStudy (anonymous):

no - do you have a calculator there ?

OpenStudy (darkbluechocobo):

1 and yes

OpenStudy (anonymous):

what do you mean , 1 ?

OpenStudy (darkbluechocobo):

i put in cos(0) and i got 1

OpenStudy (anonymous):

ah but i don't want the cosine of zero, i want the angle that has its cosine equal to zero

OpenStudy (anonymous):

okay, i'll try a slightly different approach

OpenStudy (anonymous):

sin of 0 is 0, right ?

OpenStudy (darkbluechocobo):

Yes

OpenStudy (anonymous):

now get cosine 90 off your calculator

OpenStudy (darkbluechocobo):

-.45

OpenStudy (anonymous):

????????????????

OpenStudy (darkbluechocobo):

D: i put in cos 90 ><

OpenStudy (anonymous):

can you just get your calculator to work out the cosine of 90 degrees ?

OpenStudy (darkbluechocobo):

Hokay.. hopefully its right this time >< is it 0

OpenStudy (anonymous):

yes - so sin(0) = cos(90)

OpenStudy (anonymous):

now, another pair work out the sin of 10 degrees for me

OpenStudy (darkbluechocobo):

would this be a decimal answer

OpenStudy (anonymous):

yes

OpenStudy (darkbluechocobo):

.17

OpenStudy (anonymous):

okay , and now work out the cosine of 80 degrees

OpenStudy (darkbluechocobo):

same thing ?

OpenStudy (anonymous):

hmm, interesting

OpenStudy (anonymous):

so sin(10) = cos(80)

OpenStudy (anonymous):

you seeing the pattern yet ?

OpenStudy (darkbluechocobo):

Yes aha you go up by 10 degrees for sin you lower 10 for cos

OpenStudy (anonymous):

right ! so in general, the two angles add up to 90 degrees

OpenStudy (darkbluechocobo):

that would be sin 20 cos 70

OpenStudy (anonymous):

you got it - so what if one angle was 35 ?

OpenStudy (darkbluechocobo):

the other would 55

OpenStudy (anonymous):

problem solved : ) you should check, work out sin35cos35 and see if it really is the same as sin55cos55

OpenStudy (anonymous):

if that's the case, then both angles must have the same range, according to the original formula

OpenStudy (darkbluechocobo):

Yes this was true

OpenStudy (darkbluechocobo):

Thank you so much for going through this with me

OpenStudy (anonymous):

you're welcome